We have to make 50 ounces of a 60% acid solution. I have a 30% acid solution and an 80% acid solution to work with. How many ounces of the 30% acid solution and the 80% acid solution do I need?
Found 2 solutions by robertb, stanbon:
Answer by robertb(5830)

0.30x + 0.80(50-x) = 0.60*50
<==> 0.30x + 40 - 0.80x = 30
==> -0.50x = -10
==> x = 20 ounces of the 30% acid solution , and 30 ounces of the 80% acid solution.

Answer by stanbon(75887)

We have to make 50 ounces of a 60% acid solution. I have a 30% acid solution and an 80% acid solution to work with. How many ounces of the 30% acid solution and the 80% acid solution do I need?
-----
Equations:
t + e = 50 ounces
30t + 80e = 60*50
----------------------------
Modify for elimination:
30t + 30e = 30*50
30t + 80e = 60*50
-------------------
Subtract and sove for "e"::
50e = 30*50
e = 30 ounces (amt. of 80% needed)
t = 50-30 = 20 ounces (amt. of 30% needed)
------------
Cheers,
Stan H.
------------